# Efficient Design for Multi-user Downlink Beamforming with Reconfigurable Intelligent Surface

### v2

@@ Abstract
This paper considers downlink multi-user transmission facilitated by a reconfigurable intelligent surface (RIS). First, focusing on the multi-group multicast beamforming scenario, we develop a fast and scalable algorithm for the joint base station (BS) and RIS beamforming optimization to minimize the transmit power subject to the user quality-of-service (QoS) constraints. By exploring the structure of this QoS problem, we show that the joint beamforming optimization can be naturally decomposed into a BS multicast beamforming QoS problem and an RIS passive multicast beamforming max-min-fair (MMF) problem. We propose an alternating multicast beamforming (AMBF) algorithm to solve the two subproblems alternatingly. For the BS QoS subproblem, we utilize the optimal multicast beamforming structure to obtain the BS beamformers efficiently. Furthermore, we reformulate the challenging RIS MMF subproblem and employ a first-order projected subgradient algorithm (PSA), which yields closed-form updates. The computational complexity of the AMBF algorithm grows linearly with the number of RIS elements and BS antennas. We further show that the AMBF approach is also an efficient method for the RIS-assisted downlink multi-user unicast beamforming problem, providing semi-closed-form updates. Next, we study the MMF problem for the RIS-assisted downlink beamforming design and propose a PSA-based fast algorithm to compute the BS and RIS beamforming solutions with closed-form updates per iteration, leading to a highly computationally efficient solution. Simulation results show the efficacy of our proposed algorithms in both performance and computational cost compared to other alternative methods.
@@ I Introduction
Reconfigurable intelligent surface (RIS) is an emerging technology that can actively control and improve the wireless propagation channel conditions and thus creates a smart reconfigurable wireless environment to enhance communications performance [2, 3]. RIS uses a planar surface consisting of passive reflective elements to control the phase shifts of the reflected wireless signal towards the desired direction with little energy usage, thus forming passive beamforming to enhance the transmission performance. It offers a range of advantages over the traditional options (e.g., base station (BS) or relay), such as no need for signal processing, low energy consumption and hardware cost, convenience for deployment, etc. Thus, RIS is a promising technology that has been explored to support a wide range of applications for next-generation wireless systems, such as coverage extension, capacity enhancement, improved localization and sensing [4, 5, 6, 7].
@@
For downlink transmissions, besides the typical dedicated unicast data transmission to each user, wireless multicasting has become increasingly popular to support content distribution and delivery in emerging applications. RIS can be explored to enhance multicast transmission performance. In conventional wireless systems, physical-layer multicast beamforming has been widely considered an efficient transmission technology for the BS to deliver common data to multiple users simultaneously. Various optimization algorithms or signal processing techniques have been developed in the literature to provide design solutions for different transmission scenarios or systems [8, 9, 10, 11, 12, 13, 14]. For delivering common data to a group, multicast beamforming performance is typically limited by the user in the group with the worst channel condition. Such limitation is dictated by the wireless prorogation environment. RIS can be utilized to break this barrier by adding an additional controllable channel path. In particular, it can form passive beamforming by controlling the phase shifts of the reflective elements to improve the user channel conditions to further enhance the overall transmission performance. Thus, it is important to explore this potential for RIS-assisted multicast beamforming design.
In conventional systems, downlink multicast beamforming is a more challenging design problem than multi-user unicast beamforming, as the former is in general an NP-hard problem [8]. This is also the case for the design of RIS-assisted downlink beamforming. RIS-assisted multicast beamforming design has recently been investigated for the single-group [15, 16] and multi-group settings [17, 18, 19, 20]. Several approaches for joint BS and RIS beamforming optimization have been proposed for either maximizing sum group rate [17, 18], or minimizing the BS transmit power subject to the user quality-of-service (QoS) target constraints, i.e., the QoS problem [19, 20]. In these works, the alternating optimization (AO) technique is commonly adopted for joint BS and RIS beamforming optimization. Depending on the specific problem formulation, smoothing techniques or SDR-based approaches have been proposed to solve the subproblems in each AO iteration. However, some limitations exist in these existing methods. For the QoS problem, existing studies are limited, and the efficacy of the proposed AO approaches in [19, 20] is uncertain as there is no guarantee of the quality of the computed solution. Also, the above existing methods generally have high computational complexity, which cannot be effectively implemented for a large number of RIS elements and BS antennas. Since RIS is beneficial only when the number of reflective elements becomes large, developing an effective design with both high performance and low computational complexity is crucial.
@@ I-A Contribution
In this paper, we consider the RIS-assisted downlink transmission design focusing on the multicast beamforming scenario. We develop our efficient fast algorithms for joint BS and RIS beamforming optimization for both the QoS problem and the max-min fair (MMF) problem, where the MMF problem refers to maximizing the weighted minimum SINR among users subject to the BS transmit power budget. Furthermore, we generalize our proposed approach and show that it can be extended to the unicast beamforming scenario, lending to a fast beamforming solution.
@@
Several unique features of our proposed algorithms distinguish our work from the existing ones. First, we propose a fast alternating multicast beamforming (AMBF) approach for the QoS problem. Different from the existing methods that directly apply AO to this problem, we explore the structure of the joint BS and RIS beamforming optimization and show that this QoS problem can be naturally broken down into two multicast beamforming subproblems, which are solved alternatingly: a BS multicast beamforming QoS problem and an RIS passive multicast beamforming MMF problem. This structure-based AMBF approach can effectively improve the BS and RIS beamforming solutions over iterations, avoiding the issue faced in the existing direct AO approaches.
@@
Next, with the two subproblems identified, we can judiciously exploit the optimal multicast beamforming structure [11] to compute a solution efficiently for the BS multicast problem, whose computational complexity does not depend on the number of BS antennas. This departs from the existing methods that are purely based on computational optimization techniques. For the challenging RIS passive multicast MMF problem, we develop our technique to reformulate the problem and adopt a first-order projected subgradient algorithm (PSA) [12] to solve it, which yields computationally cheap closed-form updates and has a convergence guarantee. Finally, we show that since our AMBF approach utilizes the structure of the problem, it is readily applicable to the downlink multi-user unicast beamforming scenario to compute the solution efficiently. Our fast algorithm is scalable, with computational complexity growing linearly with the number of RIS elements and the BS antennas.
For the RIS-assisted downlink multicast MMF problem, we first show it is an inverse problem to the earlier QoS problem. Following this, we utilize the optimal BS beamforming structure again to transform the MMF problem into an equivalent problem, which allows us to propose a PSA-based iterative algorithm to compute the BS and RIS beamforming solutions. In particular, the algorithm uses closed-form updates per iteration, leading to a highly computationally efficient solution. The proposed algorithm is also directly applicable to the general downlink RIS-assisted multi-user beamforming problem with the MMF objective.
The simulation results demonstrate the effectiveness of our proposed algorithms, in terms of the performance and computational advantage, for both the multicast and unicast scenarios compared to the alternative methods.
I-B Related Work
Multicast beamforming optimization is inherently a non-convex and NP-hard problem [8]. The earlier literature works adopted the semi-definite relaxation (SDR) approach [8, 9] to find an approximate solution. With the number of antennas increasing, the successive convex approximation (SCA) technique has become more attractive for its lower computational complexity and better performance [10]. The optimal multi-group multicast beamforming structure was recently obtained in [11]. It shows that to obtain the beamforming vectors, we only need to optimize the per-user weight, which is a much smaller optimization problem that is independent to the number of BS antennas. This structure has been exploited to develop fast and scalable algorithms for multicast beamforming in massive multiple-input multiple-output (MIMO) systems [12, 13, 14].
Several recent works have studied the RIS-assisted multicast beamforming design for the single group [15, 16] and the general multi-group settings [17, 18, 19, 20]. Joint BS and RIS beamforming optimization for maximizing sum group rate is considered in [17, 18]. In [17], the alternating optimization (AO) technique is used, and a smoothing technique is proposed along with a majorization minimization method to solve the subproblems. In [18], the smoothing technique is applied directly to the rate objective, along with a low-complexity alternating projected gradient algorithm. The QoS problem for the BS transmit power minimization is studied in [19], where AO is directly applied to the problem, resulting in a feasibility problem for the RIS passive beamforming design. The SDR-based approach is then proposed to solve the subproblems. A similar approach is also considered in [20] for an RIS-assisted symbiotic radio multicast system comprising the primary receivers and an internet-of-things receiver. However, as the feasibility problem cannot effectively search for an improved solution over iteration, there is no guarantee of the quality of this AO approach.
The unicast scenario for RIS-assisted downlink multi-user beamforming has been studied in many recent works. Among them, only a few works have studied the BS power minimization problem [5, 21]. In [5], directly applying AO to the joint beamforming optimization problem is first considered, and a low-complexity heuristic method based on the sum of channel gains is proposed. RIS-assisted non-orthogonal multiple access (NOMA) network is considered in [21], and the joint beamforming design using the direct AO approach is also adopted. Both [5] and [21] adopt semi-definite programming (SDP) or SDR to solve the corresponding alternating subproblems. The minimum rate (or SINR) maximization problem under various system configurations has been considered in [22, 23, 24]. In [22], an iterative algorithm providing an alternating descent path is proposed. The work in [23] uses asymptotic analysis and some assumptions for the channel model to provide the approximations to the optimal BS transmit beamformer. Then, a projected gradient ascent algorithm is proposed for the RIS beamformer based on the large-scale channel statistics. RIS-assisted multi-cell coordinated transmit beamforming is considered in [24], where RIS can adjust both reflecting phases and amplitudes. The AO approach is adopted, and the second-order cone programming (SOCPs) and SDR are used to solve the subproblems. An inexact AO approach with SCA is further considered to reduce the computational complexity. However, these existing approaches generally have computational complexity.
I-C Organization and Notations
@@
The rest of this paper is organized as follows. Section II presents the system model and the problem formulation for the RIS-assisted multicast beamforming QoS problem. In Section III, we develop our AMBF approach to break the QoS problem into two subproblems to solve alternatingly. In Section IV, we present the fast algorithms to solve each beamforming subproblem. In Section V, we generalize our approach to solve the multi-user unicast beamforming. The MMF problem is considered in Section VI, where we discuss the relation between the QoS and MMF problems and propose a fast first-order algorithm to solve the MMF problem. The simulation results is provided in Section VII, followed by the conclusion in Section VIII.
@@ II-A System Model
We consider an RIS-assisted downlink multi-group multicast scenario: a BS equipped with $N$ antennas multicasts messages to $G$ user groups, and an RIS consisting of $M$ passive reflective elements is deployed to assist the data transmission between the BS and the users. We assume that group $i$ consists of $K_{i}$ single-antenna users, and these users receive a common message from the BS that is independent of other groups. The total number of users is denoted by $K_{\text{tot}}=\sum_{i=1}^{G}K_{i}$. The BS controls the phases of the RIS array elements via a RIS controller to adjust the signals received at the RIS to the desired directions. We denote the set of group indices by ${\cal G}\triangleq\{1,\ldots,G\}$, the set of user indices in group $i$ by ${\cal K}_{i}\triangleq\{1,\ldots,K_{i}\}$, $i\in{\cal G}$, and the set of RIS element indices by ${\cal M}\triangleq\{1,\ldots,M\}$.
Let ${\bf H}_{\text{r}}$ denote the $M\times N$ channel matrix from the BS to the RIS, ${\bf h}^{\text{d}}_{ik}$ the $N\times 1$ channel vector from the BS to user $k$ in group $i$, and ${\bf h}^{\text{r}}_{ik}$ the $M\times 1$ channel vector from the RIS to user $k$ in group $i$, for $k\in{\cal K},i\in{\cal G}$. We use ${\bf w}_{i}$ to represent the $N\times 1$ multicast beamforming vector at the BS for group $i\in{\cal G}$. Also, ${\bf e}\triangleq[e_{1},\ldots,e_{M}]^{T}$ represents the vector containing the $M$ reflection coefficients at the RIS, where $e_{m}=e^{j\theta_{m}}$ with $\theta_{m}\in(-\pi,\pi]$ being the phase shift of element $m$. We refer to ${\bf e}$ as the RIS passive beamforming vector. The signal received at user $k$ in group $i$ is given by
$\!\!y_{ik}$ $=\sum_{j=1}^{G}{\bf w}_{j}^{H}\big{(}{\bf h}^{\text{d}}_{ik}+{\bf H}_{\text{r}}^{H}\text{diag}({\bf e}){\bf h}^{\text{r}}_{ik}\big{)}s_{j}+n_{ik}$
$={\bf w}_{i}^{H}\!\big{(}{\bf h}^{\text{d}}_{ik}\!+\!{\bf G}_{ik}{\bf e}\big{)}s_{i}+\!\sum_{j\neq i}\!{\bf w}_{j}^{H}\!\big{(}{\bf h}^{\text{d}}_{ik}\!+\!{\bf G}_{ik}{\bf e}\big{)}s_{j}+n_{ik}$ (1)
@@
$\textrm{SINR}_{ik}=\frac{|{\bf w}_{i}^{H}\big{(}{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}\big{)}|^{2}}{\sum_{j\neq i}|{\bf w}_{j}^{H}\big{(}{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}\big{)}|^{2}+\sigma^{2}},$ (2)

### v3

# Fast and Scalable Beamforming for RIS-Assisted Downlink Multi-group Multicasting

@@ Abstract
This paper considers downlink multi-group multicasting via beamforming facilitated by a reconfigurable intelligent surface (RIS). We develop a fast and scalable algorithm for the joint base station (BS) and RIS beamforming optimization to minimize the transmit power while meeting user quality-of-service (QoS) targets. By analyzing the structure of the QoS constraints, we reformulate the problem and show that the joint beamforming optimization inherently consists of a multicast beamforming QoS problem for the BS and a passive multicast beamforming max-min-fair (MMF) problem for the RIS. We propose a fast alternating multicast beamforming (AMBF) algorithm to effectively solve the two subproblems alternatingly. For the BS multicast subproblem, we utilize the optimal multicast beamforming structure to efficiently determine the BS beamformers. For the RIS multicast subproblem, we reformulate the MMF problem and apply a first-order projected subgradient algorithm (PSA), which yields simple closed-form updates. The computational complexity of the AMBF algorithm grows linearly with the number of RIS elements and BS antennas. We further consider joint BS and RIS beamforming for the weighted MMF design objective subject to the BS transmit power budget. We propose an alternating PSA (APSA) fast algorithm to compute the beamforming solutions for the BS and RIS. APSA consists only closed-form updates per iteration, yielding linear computational complexity in the number of RIS elements and BS antennas. Simulation results show the efficacy of our proposed algorithms in terms of performance and computational cost compared to alternative methods.
@@ I Introduction
Reconfigurable intelligent surface (RIS) is an emerging next-generation technology that can actively control and enhance wireless propagation channel conditions, thereby creating a smart reconfigurable wireless environment to enhance communications performance [2, 3]. RIS uses a planar surface consisting of passive reflective elements to control the phase shifts of the reflected wireless signal towards desired directions with minimal energy consumption. This forms passive beamforming, which enhances the transmission performance. RIS offers several advantages over the traditional methods (e.g., base station (BS) or relay), such as the elimination of the need for complex signal processing, low energy consumption, reduced hardware cost, and convenience for deployment, etc. The potential role of RIS in next-generation wireless networks has been investigated to support a variety of applications, including coverage extension, capacity enhancement, and improved localization and sensing [4, 5, 6, 7, 8].
@@
Existing and emerging wireless services and technologies, such as video streaming, software dispatch distribution, edge storage and computing, and distributed machine learning, are driving significant growth of downlink content and data distribution. This demand increases the need of wireless multicast strategies to support efficiency delivery of shared content. In particular, with a multi-antenna BS, physical-layer multicasting via beamforming is an efficient transmission technique for the BS to deliver common data to multiple users simultaneously, which can effectively improve spectrum and power efficiency. Numerous optimization algorithms or signal processing techniques have been developed [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. For delivering common data to a group, multicast beamforming performance is typically limited by the user in the group with the worst channel condition. This limitation arises from the wireless prorogation environment. RIS can be utilized to enhance multicast transmissions by adding an additional controllable channel path to mitigate this condition. By adjusting the phase shifts of the reflective elements, RIS enables passive beamforming that enhances the channel conditions for users, thereby improving the overall transmission performance. Therefore, it is crucial to explore effective and efficient designs for RIS-assisted multicast beamforming.
Downlink multicast beamforming presents a challenging design problem, as it is generally a non-convex NP-hard problem [9]. Earlier literature adopted the semi-definite relaxation (SDR) approach [9, 10, 11] to find approximate solutions. However, with the increasing number of BS antennas in evolved wireless systems, the successive convex approximation (SCA) technique has gained popularity due to its improved performance and lower computational complexity [12]. To further reduce the solution complexity in massive multiple-input multiple-output (MIMO) systems, various computational methods or signal processing techniques have been developed for single-cell scenarios [13, 14, 15], multi-cell cooperation [16, 17], as well as robust designs that account for imperfect channel state information [18, 19]. The optimal structure for multi-group multicast beamforming has been established recently in [20]. It indicates that to design the BS beamformers, we only need to optimize the weights among users in each group. This significantly simplifies the optimization problem, as it is much smaller and independent of the number of BS antennas. This optimal structure has been leveraged to develop scalable ultra-low-complexity algorithms for massive MIMO downlink multicasting [21, 22], mixed traffic scenarios [23], and multi-cell coordination[24].
The challenge of multicast beamforming design extends to RIS-assisted downlink multicasting. RIS-assisted multicast beamforming design has been investigated for the single-group [25, 26] and multi-group settings [27, 28, 29, 30]. Several approaches have been proposed for designing joint BS and RIS beamforming [27, 28, 29, 30]. The alternating optimization (AO) technique is commonly adopted for joint BS and RIS beamforming optimization. Joint BS and RIS beamforming optimization for maximizing the sum group rate is considered in [27, 28]. In [27], a smoothing technique is proposed along with a majorization minimization method to solve the subproblems in each AO iteration. In [28], the smoothing technique is applied directly to the rate objective, along with a low-complexity alternating projected gradient algorithm. However, while maximizing the sum group rate can enhance overall data rate, it does not guarantee fairness among different groups. As a result, some groups may experience significant performance degradation, leading to an unacceptable user experience. To tackle the QoS issue, the QoS problem that minimizes the BS transmit power while meeting user QoS targets is studied in [29], where AO is directly applied to the optimization problem, resulting in a feasibility subproblem for the RIS passive beamforming design. The SDR-based approach is then proposed to solve the non-convex subproblems in each AO iteration. A similar approach is also considered in [30] for an RIS-assisted symbiotic radio multicast system comprising the primary receivers and an internet-of-things receiver. However, solving the feasibility problem does not ensure the quality of the solution for the RIS passive beamforming subproblem, raising questions about the effectiveness of this direct AO approach. Additionally, the methods mentioned above generally entail high computational complexity, which hinders their practical implementation. Most of these studies have only considered a relatively small number of RIS elements for performance evaluation. Since RIS systems typically require hundreds of elements to be beneficial, it is crucial to develop a scalable design solution that balances both high performance and low computational complexity.
@@ I-A Contribution
This paper aims to provide an effective joint BS and RIS beamforming design for RIS-assisted downlink data multicasting. We develop fast and scalable algorithmic solutions for two multicast beamforming formulations: the QoS problem, which minimizes the BS transmit power and the max-min fair (MMF) problem, which maximizes the weighted minimum signal-to-interference-and-noise ratio (SINR) among users, subject to the BS transmit power limit.
@@
We propose a fast alternating multicast beamforming (AMBF) approach for solving the QoS problem. This approach differs from existing methods that directly apply the AO technique to the joint optimization problem. Instead, we explore the structure of the constraints and reformulate the problem to reveal that the QoS problem inherently consists of two separate multicast beamforming optimization subproblems: a conventional multicast beamforming QoS problem for the BS and a passive multicast beamforming MMF problem for the RIS. The proposed AMBF approach solves these two subproblems alternatingly, effectively improving the BS and RIS beamforming solutions over iterations.
@@
We design fast algorithms to effectively solve the two subproblems. For the BS multicast problem, we utilize the optimal multicast beamforming structure [20] to compute a solution efficiently, where the computational complexity is independent of the number of BS antennas. This approach contrasts with existing methods for RIS-assisted beamforming, which rely on computational optimization techniques. For the more challenging RIS passive multicast MMF problem, we reformulate the problem to employ a first-order projected subgradient algorithm (PSA), which yields computationally inexpensive closed-form updates and has a guaranteed convergence. In the special case of the multi-user unicast beamforming scenario, the optimal BS beamformer can be computed exactly, and our AMBF approach requires only closed-form solutions or updates for both subproblems. Our AMBF algorithm is scalable, with computational complexity increasing linearly with the number of RIS elements and BS antennas.
For the MMF problem, we show that it is an inverse problem to the earlier QoS problem. We leverage the optimal BS beamforming structure to transform the MMF problem into an equivalent lower-dimensional problem and propose an alternating PSA (APSA) algorithm that efficiently computes the BS and RIS beamforming solutions. Our APSA algorithm employs closed-form updates for both BS and RIS beamformers, maintaining linear computational complexity with respect to (w.r.t.) the number of RIS elements and BS antennas.
Simulation results show the effectiveness of our proposed algorithms, demonstrating superior performance compared to existing AO-based algorithms and comparative to SCA-based joint optimization methods, while achieving significant reduction in computational time.
I-B Organization and Notations
@@
The rest of this paper is organized as follows. Section II presents the system model and the problem formulation for the RIS-assisted multicast beamforming QoS problem. In Section III, we develop our AMBF approach to break the QoS problem into two subproblems to solve alternatingly. In Section IV, we present the fast algorithms to solve each beamforming subproblem. The MMF problem is considered in Section V, where we discuss the relation between the QoS and MMF problems and propose the APSA algorithm to solve the MMF problem efficiently. The simulation results is provided in Section VI, followed by the conclusion in Section VII.
@@ II-A System Model
We consider an RIS-assisted downlink multicast scenario: a BS equipped with $N$ antennas multicasts messages to $G$ user groups, and an RIS consisting of $M$ passive reflective elements is deployed to assist the data transmission between the BS and the users, as shown in Fig. 1. We assume that group $i$ consists of $K_{i}$ single-antenna users, and the BS sends a common message to this group of users that is independent of other groups. The total number of users is denoted by $K_{\text{tot}}=\sum_{i=1}^{G}K_{i}$. The BS controls the phases of the RIS array elements via a RIS controller to adjust the signals received at the RIS to the desired directions. We denote the set of group indices by ${\cal G}\triangleq\{1,\ldots,G\}$, the set of user indices in group $i$ by ${\cal K}_{i}\triangleq\{1,\ldots,K_{i}\}$, $i\in{\cal G}$, and the set of RIS element indices by ${\cal M}\triangleq\{1,\ldots,M\}$.
[…]
Figure 1: An example of RIS-assisted downlink multi-group multicasting.
Our design considers a general scenario where both the direct path and the RIS path are available between the BS and the users, [^1] In a rich scattering environment, although there may not be a line-of-sight path available, the non-line-of-sight paths typically still exist between the BS and a user. which includes the special case that only the RIS path is available. Let ${\bf H}_{\text{r}}$ denote the $M\times N$ channel matrix from the BS to the RIS, ${\bf h}^{\text{d}}_{ik}$ the $N\times 1$ channel vector from the BS to user $k$ in group $i$, and ${\bf h}^{\text{r}}_{ik}$ the $M\times 1$ channel vector from the RIS to user $k$ in group $i$, for $k\in{\cal K},i\in{\cal G}$. We use ${\bf w}_{i}$ to represent the $N\times 1$ multicast beamforming vector at the BS for group $i\in{\cal G}$. Also, ${\bf e}\triangleq[e_{1},\ldots,e_{M}]^{T}$ represents the vector containing the $M$ reflection coefficients at the RIS, where $e_{m}=e^{j\theta_{m}}$ with $\theta_{m}\in(-\pi,\pi]$ being the phase shift of element $m$. We refer to ${\bf e}$ as the RIS passive beamforming vector. The signal received at user $k$ in group $i$ is given by
$y_{ik}$ $=\!\big{(}{\bf h}^{\text{d}}_{ik}+{\bf H}_{\text{r}}^{H}\text{diag}({\bf e}){\bf h}^{\text{r}}_{ik}\big{)}^{\!H}\!\sum_{j=1}^{G}{\bf w}_{j}s_{j}\!+n_{ik}$
$=\!\big{(}{\bf h}^{\text{d}}_{ik}\!+\!{\bf G}_{ik}{\bf e}\big{)}^{\!H}{\bf w}_{i}s_{i}+\!\sum_{j\neq i}\!\big{(}{\bf h}^{\text{d}}_{ik}\!+\!{\bf G}_{ik}{\bf e}\big{)}^{\!H}{\bf w}_{j}s_{j}\!+\!n_{ik}$
@@
$\textrm{SINR}_{ik}=\frac{|\big{(}{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}\big{)}^{H}{\bf w}_{i}|^{2}}{\sum_{j\neq i}|\big{(}{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}\big{)}^{H}{\bf w}_{j}|^{2}+\sigma^{2}},$ (1)